# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v2

@@ 4.2. Physical Optimization
Given an optimized logical plan, the physical optimizer considers various implementations of the plan and returns the optimal one. We begin with a discussion of different approaches to constraint enforcement, and then outline ideas for optimized plan selection.
@@ 4.2.1. Enforcement Implementations
Constraints can be enforced in multiple ways, typically categorized as reactive or proactive, based on whether enforcement occurs after or during response generation, respectively. While the DPS can enforce all constraints reactively, only certain ones can be enforced proactively.
@@
Proactive Enforcement. Proactively enforcing constraints can lower costs significantly. For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, identifying and masking invalid tokens to ensure compliance by construction. Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token validation. However, constrained decoding is not limited to grammar-based approaches. For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints. Yet, open source libraries currently do not offer this functionality.
An open challenge is to identify which other SICs can be enforced via constrained decoding. Moreover, developing a framework to efficiently apply various constraints on different output attributes also remains unresolved. Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024). Nonetheless, it provides a foundation for future exploration. Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions. This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
Besides deterministic constrained decoding, the DPS can proactively enforce stochastic constraints by asynchronously detecting violations in partial streamed outputs. This approach is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint. For instance, the DPS can proactively enforce a grounding constraint on an abstractive operator by asynchronously validating generated summary sentences with a fact-checking model (Tang et al., 2024). Upon a violation, the operator can backtrack to the decoding step right before the violation. Alternative techniques like re-prompting with partial feedback are also worth considering to address this problem.
@@ 4.2.2. Optimized Plan Selection
For a fixed set of constraints and their implementations, the physical optimizer can adapt existing techniques (Liu et al., 2025b; Russo et al., 2025) to select an optimal query plan defined by user-specified cost or reliability (a.k.a. quality) thresholds (omitted from the query interface for brevity). Specifically, rather than measuring the reliability of an operator’s implementation in terms of validation data or the best implementation from prior beliefs (e.g., a “champion model”), reliability is instead defined by the proportion of output tuples that satisfy all of the operator’s constraints (after necessary retries). However, the challenge is selecting each constraint’s implementation to begin with.
As discussed in Section 4.2.1, there are many ways to detect a constraint violation, each with different cost-reliability trade-offs. Formally, the cost of a constraint’s implementation is the overhead of detecting a violation in a single output. The reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation. A deterministic constraint is completely reliable (i.e., 100% precision and recall), while a stochastic constraint is not completely reliable, due to invoking models such as LLMs.
To choose the best constraint implementation, we must estimate costs and reliability for each option. While cost estimation is often simple and can be done during runtime with sampled data, it might not always be necessary. For instance, proactively enforcing a deterministic constraint through constrained decoding is usually quicker than doing it reactively.
Assessing the reliability of stochastic constraints is more challenging. Unlike deterministic constraints that are always reliable, the reliability of stochastic constraints depends on (1) how they are implemented, (2) the output tuples from upstream operators in the query plan, and (3) the input tuples to the query. Moreover, obtaining robust reliability estimates for constraint implementations can be difficult, as they often need ground-truth labels from human annotators. We present several strategies to address this issue.
One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime. This approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates. Since metrics are averaged across diverse query contexts, they do not effectively reflect interactions with upstream operators. Furthermore, data and workload drift may make these estimates obsolete. It is also generally costly for organizations to annotate execution traces.
@@
A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., expected confidence) of a constraint’s implementation at runtime based on sampled data. Rather than defining precision and recall thresholds, users specify confidence thresholds instead. Although many conventional classifiers inherently provide confidence scores, they can be derived from LLMs by applying softmax to the “True” and “False” token logits. The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability. Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks. Though, further research is needed to obtain highly-calibrated constraints.
@@ 5. Observability
After a user submits a query to the DPS, the observability stack enables the user to understand the performance of their query’s operators and constraints. The stack has three components. The observability interface allows the user to view relevant metrics (e.g., cost, reliability) and inspect query results. The observability store maintains metrics and outputs from each stage in the query. Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint. While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread in both academia and industry (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), their reliability remains uncertain. As a result, labeling interfaces remain essential. A key open challenge lies in ensuring that the observability store can support the proposed interfaces efficiently at scale and that these interfaces are effectively designed. We detail each component of the stack below.
@@
Observability Interface. Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity. However, to observe the reliability for stochastic constraints, the user must submit a labeling request for their query. Once the labeling is complete, the interface displays the precision and recall of each constraint. To avoid confusion, the interface indicates whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well. The observability interface also allows the user to inspect individual output tuples. A tuple is annotated if any upstream operator violated a constraint but CONTINUEed to propagate the error. The user can view the tuple’s lineage for more fine-grained debugging.
@@
Observability Store. The observability store is modeled as additional relations in the DPS. It maintains the lineage of each query, enabling systematic debugging and performance monitoring. To obtain human annotations, the DPS can draw concepts from crowd-sourced databases (Franklin et al., 2011; Marcus et al., 2011; Park et al., 2012), such as crowd-sourced attributes and user interface generation; however, the DPS need not implement these ideas in their full generality. For example, suppose the DPS stores a relation where each tuple represents an invocation of a constraint. The relation includes attributes for the constraint’s input and output (i.e., predicted label) among others (e.g., query ID). In particular, the relation also has a crowd-sourced attribute for the true label (initially null). When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample of the relation) the precision and recall of each constraint for the given query ID. The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
@@
Labeling Interface. A key aspect of crowd-sourced databases is their ability to automatically create user interfaces for crowd workers based on declarative queries. Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs. The annotator’s role is to determine if the inputs meet the constraint. To avoid bias, the original predicted label should be hidden.
@@ 6. Enterprise-Wide Constraints
Auto-Guardrails and Recommendations. Although declaring SICs is relatively straightforward, determining the applicable constraints is more difficult. This requires anticipating the potential failure cases for each semantic operator in the query. Recent systems have been developed to assist users in formulating constraints (Shankar et al., 2024b, c), but these approaches typically generate suggestions based solely on the contents of the user’s query. In enterprise settings, constraint recommendations can be greatly enriched. Organizations typically enforce domain-specific rules that necessitate certain constraints for all pertinent queries. For instance, a hospital might require excluding PII in queries unless essential for analysis. Moreover, different teams within the same organization might use similar queries with shared constraints. Thus, the constraint store needs to automatically capture and manage common constraints organization-wide, facilitating their reuse and recommendation in related queries. We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries. This strategy boosts the efficiency of similarity searches (Malkov and Yashunin, 2020) for identifying relevant constraints, which can then be refined by an LLM to match the user’s context.
Conflicts. Conflicts can arise when queries contain multiple SICs (e.g., from accumulating shared constraints across an organization’s workloads). For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy both constraints without redaction or abstraction. Likewise, domain constraints over disjoint value sets can lead to inevitable failures. The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings. Users can address conflicts by adjusting priorities (e.g., making one constraint as soft), relaxing scope, or merging constraints as needed. This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
@@ 7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases.
ML assertions were introduced to monitor and enhance model performance by applying UDFs to model outputs to find potential errors (Kang et al., 2020). Building upon this idea, recent developments broaden assertions for LLMs by employing LLMs as judges and ML classifiers for both offline evaluation (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024). Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts). These tools complement SICs and can suggest declarative constraints to boost enforcement reliability and efficiency. Our vision for enterprise constraints also draws inspiration from these efforts. DSPy Assertions enable users to define assertions in LLM pipelines using UDFs, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024). While SICs use iterative refinement, they do not currently include prompt optimization capabilities, although they can be added. To the best of our knowledge, no framework provides a fully declarative method for optimized constraint enforcement.
A related area of work is constrained decoding, which proactively enforces constraints so that the LLM’s response adheres to them by construction (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). Current frameworks focus on grammar-based approaches; however, similar to earlier work (Beurer-Kellner et al., 2023), we view constrained decoding as a more general approach that is not limited to expressing constraints as grammars.
Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a). They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. However, their approach to validation is largely imperative. Palimpzest does not support constraints, but they allow users to specify quality thresholds to select optimized physical plans (Liu et al., 2025b; Russo et al., 2025).
@@ 8. Conclusions
SICs offer a unified, declarative foundation for improving the reliability of AI-augmented data processing. By generalizing traditional integrity constraints to the semantic behaviors of LLM-augmented operators, SICs bridge the gap between declarative query processing and the uncertainty of generative models. Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse. Through this framework, we outline a principled path toward building auditable, trustworthy AI-augmented systems capable of supporting mission-critical workflows. SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.

### v3

@@ 4.2. Physical Optimization
Given an optimized logical plan, the physical optimizer considers different implementations of the plan and returns the optimal one. We first discuss approaches to constraint enforcement and then outline ideas for optimized plan selection.
@@ 4.2.1. Enforcement Implementations
Constraints are enforced reactively or proactively, based on whether enforcement occurs after or during response generation. While all constraints can be enforced reactively, only some can be enforced proactively.
@@
Proactive Enforcement. Proactively enforcing constraints can lower costs significantly. For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, masking invalid tokens to ensure compliance by construction. Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token masking. However, constrained decoding is not limited to grammar-based approaches. For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints for extractive operators. Yet, we are not aware of any existing library that offers this functionality.
An open challenge is to identify which other SICs can be enforced via constrained decoding. Moreover, developing a framework to apply efficiently various constraints on different output attributes also remains unresolved. Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024). Nonetheless, it provides a foundation for future investigation. Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions. This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
Besides deterministic constrained decoding, the DPS can proactively enforce some stochastic constraints on partial outputs. For instance, the DPS can enforce a grounding constraint on an abstractive operator by invoking a fact-checking model (Tang et al., 2024) after each sentence is generated. While recent research (Loula et al., 2025) explores these ideas, they do not focus on efficient optimizations, such as asynchronously detecting violations in partial streamed outputs. This optimization is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint. Upon a violation, the operator can backtrack to the decoding step right before the violation; though, alternative techniques like re-prompting with partial feedback are also worth considering.
@@ 4.2.2. Optimized Plan Selection
For a fixed set of constraints and their implementations, the physical optimizer can adapt existing techniques (Liu et al., 2025b; Russo et al., 2025) to select an optimal query plan defined by user-specified cost or reliability (a.k.a. quality) thresholds (omitted from the query interface for brevity). Specifically, rather than measuring the reliability of an operator’s implementation in terms of an oracle, reliability is instead defined by adherence to all of the operator’s constraints (after necessary retries). The challenge is selecting each constraint’s implementation in the first place.
As illustrated in Section 4.2.1, each constraint’s implementation has unique cost-reliability trade-offs, which allow users to specify additional thresholds to control them. While existing cost estimation approaches can be extended to account for each operator’s constraints, estimating the reliability of the constraints is a more difficult task. Formally, the reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation. A deterministic constraint is completely reliable (i.e., 100% precision and recall) and requires no estimation, while a stochastic constraint is not completely reliable due to invoking models. Estimating the reliability of stochastic constraint implementations is challenging. First, such estimates require ground-truth labels from human annotators. Furthermore, reliability is a function of both the implementation and its input, and this input, in turn, depends on the upstream query plan and the underlying data. We present some initial ideas for constraint reliability estimation.
One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime. However, this approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates. Since metrics are aggregated across diverse query contexts, they do not effectively reflect interactions with upstream operators. Furthermore, data and workload drift may make these estimates obsolete. It is also costly for individual organizations to annotate execution traces.
@@
A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., confidence) of a constraint’s implementation at runtime based on sampled data. Rather than defining precision and recall thresholds, users specify confidence thresholds instead. While many conventional classifiers inherently provide confidence scores, they can also be derived from LLMs by applying softmax to the “True” and “False” token logits. The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability. Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks. Though, further research is needed to obtain highly-calibrated constraints across varying domains.
@@ 5. Observability
After a user submits a query to the DPS, the observability stack enables the user to understand the performance of their query’s operators and constraints. The stack has three components. The observability interface allows the user to view relevant metrics (e.g., cost, reliability) and inspect query results. The observability store maintains metrics and outputs from each stage in the query. Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint. While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), they are not entirely reliable. As a result, labeling interfaces remain essential. We detail each component of the stack below.
@@
Observability Interface. Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity. However, to observe the reliability of stochastic constraints, the user must submit a labeling request for their query. Once the labeling is complete, the interface displays the precision and recall of each constraint. The interface can also indicate whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well. Moreover, the observability interface allows the user to inspect individual output tuples. A tuple is flagged if any upstream operator violated a constraint but CONTINUEed to propagate the error. The user can also view the tuple’s lineage for more fine-grained debugging.
@@
Observability Store. The observability store is modeled as additional relations in the DPS. It maintains the lineage of each query, enabling systematic debugging and performance monitoring. To obtain human annotations, the DPS can draw concepts from crowd-sourced databases (Franklin et al., 2011; Marcus et al., 2011; Park et al., 2012), such as crowd-sourced attributes and user interface generation; however, the DPS need not implement these ideas in their full generality. For example, suppose the DPS stores a relation where each tuple represents an invocation of a constraint. The relation includes attributes for the constraint’s input and output (i.e., predicted label) among others (e.g., query ID). In particular, the relation also has a crowd-sourced attribute for the true label (initially null). When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample) the precision and recall of each constraint for the given query ID. The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
@@
Labeling Interface. A key aspect of crowd-sourced databases is their ability to create user interfaces automatically for crowd workers based on declarative queries. Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs. The annotator’s role is to determine if the inputs satisfy the constraint.
@@ 6. Enterprise-Wide Constraints
Recommendations. Although declaring SICs is relatively simple, determining the applicable constraints is more difficult. This requires anticipating the potential failure cases for each semantic operator in the query. Recent systems have been developed to assist users in formulating constraints (Shankar et al., 2024b, c), but these approaches typically generate suggestions based solely on the contents of the user’s query. In enterprise settings, constraint recommendations can be greatly enriched. Organizations typically enforce domain-specific rules that necessitate certain constraints for all pertinent queries. For instance, a hospital might require excluding PII in queries unless essential for analysis. Moreover, different teams within the same organization might use similar queries with shared constraints. Thus, the constraint store needs to capture and manage common constraints automatically organization-wide, facilitating their reuse and recommendation in related queries. We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries. This approach enables similarity searches to identify relevant constraints, which can then be refined by a reranker and LLM to match the user’s context.
Conflicts. Conflicts can arise when queries contain multiple SICs, such as from accumulating shared constraints across an organization’s workloads. For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy the constraints without redaction or abstraction. Likewise, domain constraints over disjoint value sets can lead to inevitable failures. The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings. Users can resolve conflicts by adjusting priorities (e.g., making one constraint as soft) or relaxing scope as needed. This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
@@ 7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases. ML assertions were introduced to monitor and enhance model performance by applying arbitrary functions to model outputs to find potential errors (Kang et al., 2020). Recent developments broaden assertions to LLMs by employing LLM judges and ML classifiers for offline evaluations (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024). Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts). These tools complement SICs and can instead suggest declarative constraints to boost enforcement reliability and efficiency. Our vision for enterprise constraints also draws inspiration from these efforts. DSPy Assertions enable users to implement assertions in LLM pipelines as functions, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024). While SICs use iterative refinement, they do not currently support prompt optimization capabilities, although they can be added. A related area of work is controlled generation, which encompasses constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) and model-based constraints (Loula et al., 2025). To the best of our knowledge, there is no unified framework that offers declarative and optimized constraint enforcement.
Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025; Agarwal and Huang, 2025; Madden et al., 2024). So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a). They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. However, their approach to validation is largely imperative. LOTUS (Patel et al., 2024) and Palimpzest (Liu et al., 2025b; Russo et al., 2025) do not support constraints, but they allow users to specify quality thresholds relative to oracles.
@@ 8. Conclusions
SICs offer a unified, declarative foundation for improving the reliability of AI-augmented data processing. By generalizing traditional integrity constraints to the semantic behaviors of LLM-augmented operators, SICs bridge the gap between declarative query processing and the uncertainty of generative models. Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse. Through this framework, we outline a principled path toward building trustworthy AI-augmented systems capable of supporting mission-critical workflows. SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.